OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (1, -3) and (3, -5). -1 -1 1 0 one half

OpenStudy (anonymous):

Please help me with this question!!! :)

OpenStudy (texaschic101):

we will need to use the slope formula.. slope = (y2 - y1) / (x2 - x1) (1,-3) x1 = 1 and y1 = -3 (3,-5) x2 = 3 and y2 = -5 now we sub slope = (-5- (-3) / (3 - 1) slope = (-5 + 3) / 2 slope = -2/2 slope = -1 any questions ?

OpenStudy (anonymous):

No thank you:) Sorry that I didn't reply back until today. :) @texaschic101

OpenStudy (texaschic101):

no problem....sometimes a person just gets too busy....I can relate to that :)

OpenStudy (anonymous):

Would you mind helping me with a couple others? I am absolutely horrible with math... Algebra is a total nightmare for me. >_< @texaschic101